OpenStudy (anonymous):

im given a table of x and y points and i need to find the logarithmic equation that relates the two. how do i even start?

OpenStudy (anonymous):

Okay, well you just need to find the base then. What is the base of the logarithm? So basically if b is your base, x is your first column and y is your second, then b^y=x.

OpenStudy (anonymous):

So if you raise both columns to the (1/y) power, then you get b ^ 1 = x ^(1/y)

OpenStudy (anonymous):

General form for a logarithmic equation: \[y = log_bx\] So the first interesting point is where y = 1. Because when y = 1, what will your x be?

OpenStudy (anonymous):

What are the first few figures of x and y that u have?

OpenStudy (anonymous):

x=1,x=2,x=3,x=4,x=5,x=6 when y=1, y=1.189, y=1.316, y=1.414, y=1.495, y=1.565

OpenStudy (anonymous):

Everybody seems to have disappeared. The difficulty here is that there is not a logarithmic relation between x = 1 and y = 1 and I am trying to think what else the question could mean.

OpenStudy (anonymous):

well thats a formula in my notes that says lny=mlnx

OpenStudy (anonymous):

ln y = m ln x ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It should be \[y = m(ln x)\] And you're trying to find m. So if you have any x,y pair you can solve for m.

OpenStudy (anonymous):

thats what i did. im not really sure how to check it though

OpenStudy (anonymous):

So you thought to ask us:-)

OpenStudy (anonymous):

well when i first posted it i hadnt found that formula

OpenStudy (anonymous):

Actually I'm not sure how that can be right, cause we still have the whole problem with (1,1). If x is 1, then we have: \[1 = m(ln\ 1)\] Which has no solutions for m since the ln(1) = 0

OpenStudy (anonymous):

Do you have the actual question handy?

OpenStudy (anonymous):

find the logarthmic equation that relates y and x. then theres the table of points i gave earlyer

OpenStudy (anonymous):

x=1,x=2,x=3,x=4,x=5,x=6 when y=1, y=1.189, y=1.316, y=1.414, y=1.495, y=1.565

OpenStudy (anonymous):

I wonder could it be log something +1....

myininaya (myininaya):

how can we? shouldn't most of those y's be irrational?

OpenStudy (anonymous):

I think they are just approximations...

myininaya (myininaya):

i dont like approximations when we are trying to find an equation to relate x to y

OpenStudy (anonymous):

Welcome to the real world:-)

myininaya (myininaya):

we can only find an approximate equation theen

OpenStudy (anonymous):

If it's ln (or log) +1 that gets us out of the 0 problem..

OpenStudy (anonymous):

The numbers roughly behave like a log o some sort...

myininaya (myininaya):

y=x^(1/4)

myininaya (myininaya):

thats the approximate relation

myininaya (myininaya):

i remember 2^(1/2)=1.41 blah blah i seen that number for x=4 so i was like 4^(1/4)=1.41 blah blah does this pattern work for the rest and then i was like oh yeah it does!

myininaya (myininaya):

why we were talking about log?

OpenStudy (anonymous):

That's the question, find the logarithmic relation, not just the relation. And the (1,1) stumped us. It looks a bit like lnx shuftied up a bit.

OpenStudy (anonymous):

i need to find log equations

OpenStudy (anonymous):

so tthe relations is y=x^1/4?

myininaya (myininaya):

yes

OpenStudy (anonymous):

so whats the log equation ?

myininaya (myininaya):

no log

OpenStudy (anonymous):

Maybe it's just log +1.

OpenStudy (anonymous):

So I suppose then that: \[ln y = \frac{1}{4}lnx\] Which is the form he was asking for originally.. Just seems like a silly way to write that relation.

myininaya (myininaya):

well yeah you can take log of both sides of y=x^(1/4)

OpenStudy (anonymous):

Then we have 1 , 1.3, 1.47

OpenStudy (anonymous):

But then it goes up too fast....

myininaya (myininaya):

what are you talking about?

OpenStudy (anonymous):

If you take log (ordinary log not ln) +1 as the relation..

myininaya (myininaya):

y=x^{1/4} 1^{1/4}=1 2^{1/4}=1.189207115 3^{1/4}=1.316074013 4^{1/4}=1.414213562 5^{1/4}=1.495349781 6^{1/4}=1.56508458 honestly i have no clue why we are talking about log?

OpenStudy (anonymous):

The question asks for the logarithmic relation between the 2 lists.

OpenStudy (anonymous):

So the first problem is (1,1) as was discussed at the beginning...

OpenStudy (anonymous):

then \[\ln(y)=\frac{1}{4}\ln(x)\]

myininaya (myininaya):

yeah but whats the point in saying that

OpenStudy (anonymous):

none whatsoever !

OpenStudy (anonymous):

I guess...

OpenStudy (anonymous):

thats what the question asks for

myininaya (myininaya):

why would we make the equation more ugly just because it ask for a log form

OpenStudy (anonymous):

some math teacher asked the question and i guess that is what they want. just a guess though because i lost most of my mind reading powers when they got shot off in the war

OpenStudy (anonymous):

I suspect it's a very practical class...

myininaya (myininaya):

satellite am i suppose to click on that virus?

OpenStudy (anonymous):

I have figured out when he puts that up:-)

myininaya (myininaya):

im not sure who that dude is

OpenStudy (anonymous):

"I hold in my hand the last envelope"

OpenStudy (anonymous):

Yes, I think we have dealt with this question, after much ado, as they say.

OpenStudy (anonymous):

Happy, Swiker?

OpenStudy (anonymous):

yes thankyou